OpenStudy (anonymous):

The rest mass energies of a proton and a neutron are 938.3 MeV and 939.6 MeV, respectively. What is the difference in their rest masses in kilograms?

OpenStudy (anonymous):

\[E = mc^2\] is the equation for relationship between rest mass-energy. So if you take the rest energy for each and change it to joules \[938.3*10^6 * 1.6*10^-19J = 1.50128*10^-10J\] and \[939.6 * 10^6 * 1.6*10^-19J = 1.50336*10^-10J\] then the m for both is: \[1.50128*10^-10J/(3.0*10*-8m/s)^2 = 1.668*10^-27kg\] and \[1.50336*10^-10J/(3.0*10^8m/s)^2 = 1.6704*10^-27kg\] so their difference in mass is: \[(1.6704-1.668)*10^-27kg = 2.4*10^-30kg\]